a half of what John's age was 4 years ago is equal to one third of what it will be in 5 years time.How old is John
Answer by xcentaur(357)

let his age be x
then we get,


1/2(x-4)=1/3(x+5)


cross multiply,
3(x-4)=2(x+5)


3x-12=2x+10
3x-2x=12+10
x=22


So his age right now is 22 years


-Prabhat